Grade 11Physical Chemistry

Calculate the no. Of hydrogen atoms in a water droplets 1 cm radius at stp

Somesh
8 Years agoGrade 11

1 Answer

Arun
8 Years ago

Dear student!
Radius of drop = 1mm= 10-1cm

density = 1gm/cm3  

So, Volume of the drop = 4/3π r3 = 4/3 x 22/7 x(10-1)3 =88/21 x10-3 =  4.19 x10-3 ml

Now, mass = density x  volume =1x 4.19 x10-3 =4.19 x10-3 gm

Since 1 mole of water contains = 18 gm = 6.022 x 1023 molecules of water

So, 4.19 x10-3  gm water will have =  6.022 x 1023 /18 x4.19 x10-3 =  1.4 x 1020 molecules of water